November 18, 2000 - We have had another condition one storm
this week, plus one of the largest snowfalls seen in recent history.
These problems with the weather have made it difficult to travel between
McMurdo Station and Williams Field and, consequently, we have been
forced to "take a few days off".
Nevertheless, we continue to make progress in testing the ATIC
detectors and assembling the instrument. One major victory this
week was finishing the setup of the pressure vessel. As seen in
the side panel the vessel is now undergoing leaks checks. The sign
attached to the vessel was taken from the LSU Department of Physics and
Astronomy which is currently undergoing renovation. The leak
checking is currently going very well with the internal pressure holding
very close to 4 psi for the last 48 hours. Early next week we will
increase to our flight operating pressure of 8 psi and monitor the
vessel for several days.
Check out ATIC Pictures for some of our latest images. Included
are new views of the Silicon Matrix detector plus some interior views of
Scott's "Discovery" Hut taken on a recent tour. Finally, check out
the McMurdo Information page which provides some data on temperature,
snowfall and daylight amount.
The pressure vessel undergoing leak